OpenStudy (anonymous):

Water is poured into a bucket according to the rate F of t equals the quotient of 7 plus t and the quantity 2 plus t, and at the same time empties out through a hole in the bottom at the rate E of t equals the quotient of the natural log of the quantity t plus 4 and the quantity t plus 2, with both F(t) and E(t) measured in pints per minute. How much water, to the nearest pint, is in the bucket at time t = 5 minutes. You must show your setup but can use your calculator for all evaluations.

OpenStudy (anonymous):

https://gyazo.com/d2a5ac63b91f596e6febfa3e8dbd91aa

OpenStudy (anonymous):

@jim_thompson5910 @SolomonZelman

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle F(t)=\frac{t+7}{t+2} }\) \(\large\color{#000000 }{ \displaystyle E(t)=\frac{\ln(x+4)}{t+2} }\)

OpenStudy (solomonzelman):

are the functions like this?

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

The amount of water is the area under the curve.... \(\large\color{#000000 }{ \displaystyle \int \left[F(t)-E(t)\right]dt }\)

OpenStudy (anonymous):

Ohh. Ok. I got it from here

OpenStudy (anonymous):

Unless you feel like doing it for me, I don't mind :p

OpenStudy (solomonzelman):

EXAMPLE OF THE SAME CONCEPT: Same way if you were given the velocities of the cars, \(y_1(x)\) for car 1 and \(y_2(x)\) for car 2, then you will need an integral of \(y_1(x)-y_2(x)\) to find how much distance is car 1 ahead of car 2. (provided that car 1 is indeed ahead)

OpenStudy (solomonzelman):

I do like integrating, especially if someone approaches and tells me that the problem is impossible. Even more do I enjoy using elemnatry properties for harder problems, but according to the policy I can't do it. (Afterall you have to integrate too)

OpenStudy (solomonzelman):

oh, forgot to mention your limits of integration are 0 and 5.

OpenStudy (solomonzelman):

because you are looking for the water in the tank in 5 minutes.

OpenStudy (anonymous):

Ughh I have to use quotient rule right?

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle \int_0^5\left[\frac{t+7}{t+2} -\frac{\ln(x+4)}{t+2} \right]dt }\)

OpenStudy (anonymous):

Oh yeah I have a neato calculator no need for me to use quotient rule.

OpenStudy (anonymous):

9.056 final answer

OpenStudy (solomonzelman):

you can't use the quotient rule

OpenStudy (solomonzelman):

it is integration (not differentiation)

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle \int_0^5\left[\frac{t+7}{t+2} -\frac{\ln(t+4)}{t+2} \right]dt }\)

OpenStudy (anonymous):

Oh my mistake. If I wanted to manually do it I would have subtract the two and find a way to simplify it so I may integrate.

OpenStudy (anonymous):

I hope you're not integrating (unless you did it for fun) as I am aware of how it is done, no need to teach me :)

OpenStudy (solomonzelman):

I don't you think you can hande the last component ln(t+4)/(t+2). If it were ln(t+4)/(t+4) or ln(t+2)/(t+2), then you would.

OpenStudy (solomonzelman):

are you familiar with polylogarithmic functions?

OpenStudy (anonymous):

No, I am not. I don't think I need to be either at least not in my current course.